# Step-DeepResearch Technical Report

### v1

@@ 5 System Architecture
Step-DeepResearch is implemented within a basic single-agent architecture following the ReAct paradigm, reframing complex deep research tasks into a dynamic Reasoning-Action-Observation loop. Upon receiving a user query, the agent initiates a cognitive iterative process through explicit reasoning and tool invocation (Figure 4), which cycles through three core phases.
@@
Translating plans into concrete actions, the agent selects the most appropriate tool (e.g., batch_web_surfer for search, todo for tracking) to initiate precise data acquisition requests.
@@
This mechanism allows the agent to autonomously determine the exploration depth until the final report is generated, supported by the tool infrastructure detailed below.
5.1 Context Management
5.1.1 General Context Management Strategies
For general task scenarios, we integrate a hybrid context management mechanism based on Summarization and Folding. This mechanism monitors context window usage and triggers intervention when approaching a preset threshold.
Summarization Strategy.
When the threshold is triggered, the system feeds the entire current history into the model for semantic-level compression. The model generates a formalized, highly condensed natural language summary that retains key nodes and future plans, creating a dynamic checkpoint that preserves the global task state.
Folding Strategy.
This employs rule-based structured pruning, including (1) Far-end Truncation to remove early non-critical interactions, and (2) Result Folding to replace lengthy tool outputs with placeholders. This retains only the initial task description and the most recent $k$ interaction turns.
5.1.2 Context Management for DeepResearch
Unlike general tasks, deep research tasks require citing information sources in the final report. Since citation details are embedded within tool results, standard folding and summarization methods are insufficient. To address this, we implement specific Reference-Preserving processing.
Reference-Preserving Summarization.
Search tool results often contain numerous hyperlinks, including redundancy irrelevant to current reasoning and original sources essential for the final report. During context compression, this module filters out task-irrelevant content while strictly preserving original links and citation markers. A structured Reference-Preserving Summary is maintained in the compressed context, allowing the agent to accurately restore and cite sources even after multi-turn summarization.
Reference-Preserving Folding.
Search results are structured, containing links, snippets, detailed content, and citations. The “detailed content” portion often has low information density but consumes significant context. Therefore, when folding is triggered in deep research tasks, we strip the detailed content from search results, retaining only the snippets and citation information to shorten the context while preserving original references.
5.2 Deep Research Toolset
To support the long-context, high-interaction, and high-uncertainty nature of deep research, we designed a tool system centered on a Human-equivalent Workflow. The core philosophy is to match tools with the agent’s behavioral characteristics and token budget constraints, ensuring input/output formats fall within the agent’s comfort zone to reduce operational burden and hallucination while improving end-to-end throughput.
5.2.1 Authoritative Enhanced Information Acquisition
@@
Curated Authority Indexing.
We assembled a professional team to evaluate diverse online sources, selecting over 600 core authoritative sites covering official government domains, industry research institutes, international organizations, and leading academic platforms. We built independent index shards for these sites, physically and logically isolating authoritative content from low-quality SEO spam, significantly improving recall stability.
Knowledge-Dense Document Retrieval.
Targeting long-form documents such as professional literature, financial reports, and official white papers, the system utilizes a dedicated library of over 20 million high-quality items. Indexing employs Paragraph-level granularity to avoid the noise introduced by ingesting entire documents. By recalling specific semantic paragraphs, the model acquires higher-density information with lower token costs.
Authority-Aware Ranking Heuristics.
In the ranking phase, the system integrates an authority boosting factor. When semantic relevance scores are comparable, the algorithm prioritizes content from authoritative sites, ensuring that research arguments are grounded in verifiable facts. Additionally, customized search parameter interfaces allow the system to dynamically adjust search behavior based on query intent.
5.2.2 Knowledge Management & File Operations
@@
Token-Efficient Patch-based Editing.
To mitigate the token waste of “full rewrites” and the reasoning burden imposed by “Diff formats” on mid-sized models, we introduced a Patch action. The agent only needs to provide the modified fragment with minimal anchor context, while the tool performs atomic updates via fuzzy matching. In scenarios involving local polishing of long-form reports, this reduces output token costs by over 70% and significantly boosts the success rate of complex edits.
Implicit Context Management.
To eliminate the risk of context window overflow caused by excessive retrieval content, we devised a “Summary-Aware + Local Storage” two-tiered strategy. When tool results exceed a preset threshold, the system truncates the immediate return, injecting only high-density summaries into the context while persisting raw data to local temporary files. The agent performs demand-paging via file.read based on summary cues. This design effectively offloads context pressure to disk, enabling virtually infinite context support for long-horizon reasoning.
Stateful Todo Management.
To prevent goal drift in long-horizon research, the todo tool encapsulates complex CRUD operations within a unified entry point. It automatically determines “create, rewrite, or destroy” states based on the current task stack. By decoupling research progress from model weights and persisting it at the tool layer, this ensures logical consistency and goal alignment over extensive interaction trajectories.
5.2.3 Interactive Execution & Multimodal Perception
@@
Human-like Terminal Interaction (Sandbox & Tmux Integration).
All execution commands are run within a restricted MCP sandbox environment. The system integrates the tmux session management mechanism; by maintaining a persistent scrollback buffer, the agent can stably operate command-line programs with state-refreshing characteristics (such as the vim editor or real-time monitoring tools). This equips the agent with human-like interactive debugging capabilities when facing complex errors, enhancing system-level fault tolerance.
Perception-optimized Resilient Browser.
We introduce a visual redundancy elimination strategy specifically for research tasks. The system dynamically identifies visual differences by calculating the PHash (Perceptual Hash) distance between page screenshots of consecutive actions. In cases of negligible page updates, the system inhibits image-based feedback and reverts to a text-only stream. This “perception-on-demand” strategy ensures the agent maintains real-time control over page states while significantly reducing multimodal token redundancy.
Multimodal Perception Tools.
For unstructured data, we have integrated specialized modules including file_parser (document parsing), asr (audio transcription), and analyze_image (image analysis), ensuring high-quality understanding of complex research materials.
@@ 6 ADR-Bench: A Custom Deep Research Benchmark
As a product designed for complex, open-domain research tasks, DeepResearch faces significant challenges in performance evaluation, characterized by a lack of unified standards, inherent subjectivity, and high expertise thresholds. To drive model iteration scientifically and efficiently while accurately reflecting perceived user quality, we designed and implemented a dual-track evaluation framework covering both general-domain user experience and professional-domain capability assessments. This chapter systematically summarizes the core methodologies, key findings, and empirical experiences derived from our Benchmark construction and evaluation standards design. It details query composition, evaluation framework architecture, and criteria formulation, alongside the challenges and exploratory initiatives undertaken, aiming to provide a reference for the evaluation of similarly complex tasks.
@@ 6.1 Query Composition
We conducted a study of users’ everyday Deep Research usage scenarios and obtained an initial set of queries drafted by domain experts. Through multiple rounds of testing, we filtered out and revised queries that were not fully suitable for Deep Research. Based on the distribution of queries in real business settings, we categorized user queries into nine domains: Law, Computer and Information Technology, Education, Finance and Business, Science and Engineering, Social Life, Literature and Arts, Healthcare, and Politics. To construct a more rigorous evaluation framework and to better assess the performance of DeepResearch, we further divided the queries into general-domain and specialized-domain categories. We selected Law and Finance as representative specialized domains; subject-matter experts authored the test items and rubrics and verified their correctness. Each of the Law and Finance domains contains 20 expert-written tasks. The remaining seven domains were treated as general-domain categories. For these, we collected real user queries from actual business usage, manually selected high-quality, diverse, and representative items, and curated 10 tasks per domain. The distribution of queries across all domains is shown in Figure 5.
Figure 5: The query distribution of ADR-Bench
@@ 6.2 Evaluation Framework
We adopt differentiated evaluation strategies for general-domain and specialized-domain queries. For general-domain queries, user questions are typically open-ended and broad in scope, with no single standard answer. As a result, it is difficult to objectively characterize report quality through fully enumerated rubrics. Therefore, we employ a human comparative evaluation approach. Specifically, for each task, two reports generated by different models are presented to evaluators under blind-review conditions. Evaluators assign one of five categorical judgments—“Left better / Right better / Both good / Both fair / Both poor”—and score the outputs across four sub-dimensions: completeness of information, depth of content, alignment with user needs, and readability, along with written justification. Compared with absolute scoring, comparative evaluation introduces a reference anchor, effectively reducing uncertainty arising from blurred quality standards in open-domain scenarios and better reflecting users’ subjective sense of “which report feels better.” To ensure reliability and consistency, we conducted systematic training for evaluators, standardized both the overall and sub-dimension scoring criteria, and formalized the blind-review procedures. This minimizes the influence of individual preference and increases the confidence level of evaluation conclusions.
In specialized-domain evaluation, we selected Finance and Law as representative domains to assess the model’s capability to generate expert-level reports. Such tasks demand advanced domain knowledge and deeper reasoning, and relying solely on manual comparative review is insufficient to support rapid model iteration in terms of cost and efficiency. Therefore, we adopt a rubric-based automated evaluation approach: domain experts design representative, relatively constrained real-world tasks and iteratively validate and refine the rubrics through multiple rounds of cross-examination. The resulting high-quality and enumerated rubrics cover the essential elements expected in professional reports, allowing us to focus on the model’s professional competency and knowledge proficiency.
Overall, the proposed evaluation framework captures model-perceived quality differences in real operational contexts for general-domain usage, while leveraging high-quality rubrics in specialized domains to monitor professional capability. This approach balances scientific rigor, operational feasibility, and iteration efficiency. Future work will explore more scalable evaluation methodologies to address declining discriminative power and rising evaluation costs as model performance converges.
@@ 6.3 Evaluation Criteria
Based on the evaluation framework, we conducted assessments on our model as well as relevant competing products and models. In this section, we present the specific evaluation criteria and the rationale underlying their design.
For comparative evaluations in general domains, a simplified version of our evaluation criteria is shown in the table below. It is worth noting that the “accuracy” dimension is difficult to validate reliably in open-ended long-form assessment tasks such as those used for DeepResearch, and such validation would impose substantial additional labor and time costs. Therefore, we only require evaluators to verify key conclusions or data points that are directly related to the task requirements and materially affect report quality, and incorporate this requirement under the “alignment with user needs” dimension.
Moreover, because the evaluation dimensions cannot, in practice, be fully disentangled for a given task, evaluators are instructed—when assigning scores on any given dimension—to focus exclusively on the criteria relevant to that dimension without taking other dimensions into account.
@@
Dimension Description & Criteria
General Evaluation Logic Core Principle: The score is not a simple summation of sub-dimensions but depends on the query scenario. Learning/Science Queries: Prioritize clear logic and gradual deepening (Logic > Diverse Data). Decision/Comparison Queries: Prioritize hard data comparison and pros/cons analysis (Data Depth > Flowery Text). Planning/Proposal Queries: Prioritize actionable steps and creativity (Completeness > Complex Reasoning). Completeness vs. Depth: “Research” queries prioritize completeness; “Analysis” queries prioritize depth. Veto Criteria: For serious topics (News/History), factual errors or bias result in automatic rejection.
Information Completeness Focus: Breadth, perspective coverage, and avoiding key omissions. Does it cover all key aspects? (e.g., Market Size + Competition + Policy for Industry Analysis). Are obvious or significant pieces of information missing? For multi-part user prompts, are all sub-questions answered?
Content Depth Focus: Narrative depth, substantial data, evidence, and insight. Specific Data: Use of numbers/percentages (e.g., “50 billion market”) vs. qualitative descriptions (“market is large”). Deduction: Summary and logical deduction vs. simple piling of search results. Insight: Does it hit critical points and offer valuable conclusions?
Requirement Fitness Focus: Responding to explicit/implicit needs (Relevance & Correctness). Intent: Does it truly understand the background and deep intent? Constraints: Are format requests followed? (e.g., Output as “Table” or “SWOT”). Accuracy: Are key conclusions and data points fact-checked and correct?
Readability Focus: Clarity and friendliness of information organization and presentation. Visual Aids: When presenting complex information (e.g., timelines, comparisons), are appropriate charts used to help users grasp key info quickly? Structure & Layout: Are section headings, bolding, and summaries used to make the content hierarchical and scannable? Is the segmentation logical? Accessibility: Is complex content clarified with key summaries? Are examples used to help readers understand difficult concepts?
For specialized domains, we require experts to curate questions with a relatively limited solution space and certain core commonalities. The rubrics must adhere to the five core principles summarized in Table 5.
@@ 6.4 Exploratory Attempts
At the early stage of constructing the evaluation framework, we needed to segment the domains to which DeepResearch queries belong, because this affects how we construct the evaluation set and determine its distribution. In our experience, defining too few domains results in insufficient coverage of query distributions, while defining too many domains leads to frequent overlaps, where a query may simultaneously fall into multiple domains. This complicates statistical analysis of query distribution and model performance across domains, reducing statistical meaning and informational value. Ultimately, we found that for DeepResearch, defining roughly 6–12 domains is generally reasonable. It is not necessary to pursue a perfectly “correct” segmentation scheme; however, one guiding principle is that the final segmentation should effectively direct us to the right experts and evaluators for each domain.
@@
We attempted a scheme where each query produced three DeepResearch reports from different models, and human evaluators assigned 0–10 overall scores. We observed significant variance among evaluators. When restricting the scale to 0–3, the scoring failed to reflect meaningful differences between reports and could not allocate scores reasonably. For DeepResearch—where standards are non-unique and unified verification criteria are difficult—comparative evaluation yields more stable results, higher confidence, and improved efficiency.
@@ 7.1 Experimental Setup
  • LLM-based Evaluation (Research Rubrics). We employed LLM judger using a ternary grading for each criterion. Preliminary experiments revealed that the original evaluation prompts exhibited significant stochasticity and failed to accurately interpret negative criteria. To ensure the robustness and reproducibility of our results, we employed targeted prompt engineering for negative constraints in Appendix A, enforced deterministic decoding by setting the temperature to 0, and utilized an ensemble scoring mechanism based on the arithmetic mean of three independent trials per criterion.
@@
We adopt a dual-method approach to quantify economic efficiency. For LLM-based ReAct agents, we calculate costs directly based on official model API pricing and the actual volume of input and output tokens consumed on the ResearchRubrics benchmark. For commercial agent systems where direct API token statistics are unavailable, we employ a time-based extrapolation method using ReAct baselines of comparable model size. Specifically, the cost for MiniMax Agent Pro is estimated by scaling the recorded cost of our MiniMax-based ReAct agent by the ratio of their official execution times. For Kimi-Researcher and Qwen DeepResearch, we utilize the Kimi-k2-thinking-based ReAct agent as the cost baseline, scaling by the ratio of the commercial agent system’s execution time to the baseline’s execution time. Additionally, for Qwen DeepResearch, we apply a pricing adjustment factor of $1.6\times$ to reflect the official API price differential between Qwen3-Max and Kimi-K2.

### v2

@@ 5 System Architecture
Step-DeepResearch is implemented within a basic single-agent architecture following the ReAct paradigm, reframing complex Deep Research tasks into a dynamic reasoning-action-observation loop. As shown in Figure 4, upon receiving a user query, the agent initiates a cognitive iterative process through explicit reasoning and tool invocation, which cycles through three core phases.
@@
Translating plans into concrete actions, the agent selects the most appropriate tool set ( e.g., batch_web_surfer for search, todo for tracking ) to initiate precise data acquisition requests.
@@
This mechanism allows the agent to autonomously determine the exploration depth until the final report is generated. To support the characteristics of these Deep Research tasks, which involve long-horizon reasoning, strong interaction, and high uncertainty, we design a unified tool system. The system is built upon several guiding principles, including reusing human interaction logic, structuring feedback information, preserving original interaction styles, and unifying tool execution through a layered abstraction.interwoven with the agent’s execution loop. The core design philosophy consists of aspects: capability alignment, providing operational capabilities comparable to those of humans when conducting research tasks on a computer; information adaptation, emphasizing structured and complete tool outputs with high information density, while explicitly accounting for context window constraints and failure recovery; and architectural simplification, merging functionally similar tools as much as possible while preserving flexibility and expressiveness. Detailed below are the core components of this infrastructure:
Authoritative Enhanced Information Acquisition.
@@
  • Curated Authority Indexing: We assembled a professional team to evaluate diverse online sources, selecting over 600 core authoritative sites covering official government domains, industry research institutes, international organizations, and leading academic platforms. We built independent index shards for these sites, physically and logically isolating authoritative content from low-quality SEO spam, significantly improving recall stability.
  • Curated Authority Indexing: We assembled a professional team to evaluate diverse online sources, selecting over 600 core authoritative sites covering official government domains, industry research institutes, international organizations, and leading academic platforms. We built independent index shards for these sites, physically and logically isolating authoritative content from low-quality SEO spam, significantly improving recall stability.
  • Knowledge-Dense Document Retrieval: Targeting long-form documents such as professional literature, financial reports, and official white papers, the system utilizes a dedicated library of over 20 million high-quality items. Indexing employs paragraph-level granularity to avoid the noise introduced by ingesting entire documents. By recalling specific semantic paragraphs, the model acquires higher-density information with lower token costs.
  • Authority-Aware Ranking Heuristics: In the ranking phase, the system integrates an authority boosting factor. When semantic relevance scores are comparable, the algorithm prioritizes content from authoritative sites, ensuring that research arguments are grounded in verifiable facts. Additionally, customized search parameter interfaces allow the system to dynamically adjust search behavior based on query intent.
Knowledge Management & File Operations.
@@
  • Token-Efficient Patch-based Editing: To mitigate the token waste of full rewrites and the reasoning burden imposed by diff formats on mid-sized models, we introduced a patch action. The agent only needs to provide the modified fragment with minimal anchor context, while the tool performs atomic updates via fuzzy matching. In scenarios involving local polishing of long-form reports, this reduces output token costs by over 70% and significantly boosts the success rate of complex edits.
  • Implicit Context Management: To eliminate the risk of context window overflow caused by excessive retrieval content, we devised a summary aware local storage strategy. When tool results exceed a preset threshold, the system truncates the immediate return, injecting only high-density summaries into the context while persisting raw data to local temporary files. The agent performs demand-paging via file.read based on summary cues. This design effectively offloads context pressure to disk, enabling virtually infinite context support for long-horizon reasoning.
  • Stateful Todo Management: To prevent goal drift in long-horizon research, the todo tool encapsulates complex CRUD operations within a unified entry point. It automatically determines “create, rewrite, or destroy” states based on the current task stack. By decoupling research progress from model weights and persisting it at the tool layer, this ensures logical consistency and goal alignment over extensive interaction trajectories.
Interactive Execution & Multimodal Perception.
@@
  • Human-like Terminal Interaction (Sandbox & Tmux Integration): All execution commands are run within a restricted MCP sandbox environment. The system integrates the tmux session management mechanism; by maintaining a persistent scrollback buffer, the agent can stably operate command-line programs with state-refreshing characteristics (such as the vim editor or real-time monitoring tools). This equips the agent with human-like interactive debugging capabilities when facing complex errors, enhancing system-level fault tolerance.
  • Perception-optimized Resilient Browser: We introduce a visual redundancy elimination strategy specifically for research tasks. The system dynamically identifies visual differences by calculating the PHash (Perceptual Hash) distance between page screenshots of consecutive actions. In cases of negligible page updates, the system inhibits image-based feedback and reverts to a text-only stream. This strategy ensures the agent maintains real-time control over page states while significantly reducing multimodal token redundancy.
  • Multimodal Perception Tools: For unstructured data, we have integrated specialized modules including file_parser (document parsing), asr (audio transcription), and analyze_image (image analysis), ensuring high-quality understanding of complex research materials.
@@ 6 ADR-Bench: A Custom Deep Research Benchmark
As a product designed for complex, open-ended research tasks, Deep Research faces long-standing evaluation hurdles, characterized by inconsistent standards, inherent subjectivity, and high expertise thresholds. To steer model iterations scientifically and efficiently while reflecting authentic user perception, we implemented a dual-track evaluation framework covering both general-domain user experience and professional-domain capability assessments. This chapter systematically summarizes our core methodologies, key findings, and empirical experiences derived from our benchmark construction and evaluation design. It details query composition, evaluation framework architecture, and criteria formulation, alongside the challenges and exploratory initiatives undertaken, aiming to provide a reference for the evaluation of similarly complex tasks.
Figure 5: The query distribution of ADR-Bench
@@ 6.1 Query Composition
We conducted a comprehensive study of users’ everyday Deep Research usage scenarios and obtained an initial set of queries drafted by domain experts. Through multiple rounds of testing, we filtered and revised queries that were not fully suitable for Deep Research tasks. Based on the distribution of queries in real-world business settings, we categorized user queries into nine domains: Law, Computer and Information Technology, Education, Finance and Business, Science and Engineering, Social Life, Literature and Arts, Healthcare, and Politics.
To construct a more rigorous evaluation framework and better assess the performance of Deep Research, we organized these domains into two primary tracks: general and professional. Law and Finance were selected as representative professional-domain categories. For these, domain experts developed both the queries and their corresponding rubrics, which then underwent rigorous cross-validation. Each of these two domains comprises 20 specialized queries. The remaining seven domains are treated as general-domain tasks, for which we collected authentic queries from real business scenarios. From these, we manually selected high-quality, diverse, and representative samples, assigning 10 queries to each domain. The overall query distribution across these domains is illustrated in Figure 5.
@@ 6.2 Evaluation Framework
We adopt differentiated evaluation strategies for general-domain and professional-domain queries. For general-domain queries, user questions are typically open-ended and broad in scope, with no single standard answer. As a result, it is difficult to objectively characterize report quality through fully enumerated rubrics. Therefore, we employ a human comparative evaluation approach. Specifically, for each task, two reports generated by different models are presented to evaluators under blind-review conditions. Evaluators assign one of five categorical judgments—“Left better / Right better / Both good / Both fair / Both poor”—and score the outputs across four sub-dimensions: information completeness, content depth, requirement fitness, and readability, accompanied by a detailed written justification. Compared with absolute scoring, comparative evaluation introduces a reference anchor, effectively reducing uncertainty stemming from ambiguous quality standards in open-domain scenarios and more accurately capturing the user’s subjective perception of report quality. To ensure reliability and consistency, we conducted systematic training for evaluators, standardized both the overall and sub-dimension scoring criteria, and formalized the blind-review procedures. This minimizes the influence of individual preference and enhances the confidence level of evaluation conclusions.
For professional-domain evaluation, we selected Finance and Law as representative domains to assess the model’s capability in generating expert-level reports. Given that these tasks demand specialized domain knowledge and sophisticated reasoning, relying solely on manual comparative review is insufficient for rapid model iteration due to high costs and efficiency constraints. Therefore, we adopt a rubric-based automated evaluation approach: domain experts design representative and relatively constrained real-world tasks, iteratively validate and refine the rubrics through multiple rounds of expert cross-validation. These high-quality, enumerated rubrics ensure comprehensive coverage of the essential elements expected in professional reports, allowing us to rigorously measure the model’s professional competency and knowledge proficiency.
Overall, the proposed evaluation framework captures model-perceived quality differences in real operational contexts for general-domain usage, while leveraging high-quality rubrics in professional domains to monitor professional capability. This approach balances scientific rigor, operational feasibility, and iteration efficiency. Future work will explore more scalable evaluation methodologies to address declining discriminative power and rising evaluation costs as model performance converges.
@@ 6.3 Evaluation Criteria
Based on the evaluation framework, we conducted a comparative assessment of our model and relevant industry competitors. In this section, we introduce the specific evaluation criteria and the strategic rationale behind their formulation.
For comparative evaluations in general domains, a simplified version of our evaluation criteria is shown in Table LABEL:tab:eval_metrics. It is worth noting that the accuracy dimension is difficult to validate reliably in open-ended long-form assessment tasks such as those used for Deep Research, and such validation would impose substantial additional labor and time costs. Therefore, we only require evaluators to verify key conclusions or data points that are directly related to the task requirements and materially affect report quality, and incorporate this requirement under the “alignment with user needs” dimension.
Moreover, because the evaluation dimensions cannot, in practice, be fully disentangled for a given task, evaluators are instructed—when assigning scores on any given dimension—to focus exclusively on the criteria relevant to that dimension without taking other dimensions into account. For professional domains, we require experts to curate questions with a relatively limited solution space and certain core commonalities. The rubrics must adhere to the five core principles summarized in Table 5.
@@
[…]
@@ 6.4 Exploratory Attempts
At the early stage of constructing the evaluation framework, we prioritized domain segmentation for Deep Research queries, as this categorization directly informed the composition and distribution of the evaluation set. In our experience, defining too few domains results in insufficient coverage of query distributions, while defining too many domains leads to frequent overlaps, where a query may simultaneously fall into multiple domains. This complicates statistical analysis of query distribution and model performance across domains, reducing statistical meaning and informational value. Ultimately, we found that for Deep Research, defining roughly 6–12 domains was generally reasonable. It is not necessary to pursue a perfectly correct segmentation scheme; however, one guiding principle is that the final segmentation should effectively direct us to the right experts and evaluators for each domain.
@@
We attempted a scheme where each query produced three Deep Research reports from different models, and human evaluators assigned 0–10 overall scores. We observed significant variance among evaluators. When restricting the scale to 0–3, the scoring failed to reflect meaningful differences between reports and could not allocate scores reasonably. For Deep Research, where standards are non-unique and unified verification criteria are difficult, comparative evaluation yields more stable results, higher confidence, and improved efficiency.
@@ 7.1 Experimental Setup
  • LLM-based Evaluation (ResearchRubrics). We employed LLM judger using a ternary grading for each criterion. Preliminary experiments revealed that the original evaluation prompts exhibited significant stochasticity and failed to accurately interpret negative criteria. To ensure the robustness and reproducibility of our results, we employed targeted prompt engineering for negative constraints in Appendix A., enforced deterministic decoding by setting the temperature to 0, and utilized an ensemble scoring mechanism based on the arithmetic mean of three independent trials per criterion.
@@
We adopt a dual-method approach to quantify economic efficiency. For LLM-based ReAct agents, costs are calculated directly from official API pricing and actual token consumption on the ResearchRubrics benchmark. For commercial agent systems without public token statistics, we estimate costs based on the principle that execution time serves as a reasonable proxy for token workload among models of comparable size. Specifically, the cost of MiniMax Agent Pro is estimated by scaling the recorded cost of the MiniMax-M2-based ReAct agent by their execution time ratio. For Kimi-Researcher and Qwen DeepResearch, we utilize the Kimi-k2-thinking-based ReAct agent to establish a baseline, scaling its cost by execution time ratios, given that Qwen3-Max and Kimi-k2 possess similar model parameters. We further adjust the Qwen DeepResearch cost by a factor of $1.6\times$ to reflect the actual API price difference between Qwen3-Max and Kimi-k2.
7.2 Evaluation Results